OpenStudy (anonymous):

Simplify: (sin2a)/sina = (cos2a)/cosa

OpenStudy (anonymous):

2sin(a)cos(a)/sin(a) = cos(2a)/cos(a) 2cos(a) = cos(2a)/cos(a) \[2\cos^2(a) = \cos(2a)\] \[2\cos^2(a) = \cos^2(a) - \sin^2(a)\] \[\cos^2(a) = -\sin^2(a)\] 1 = 0 No solution!

OpenStudy (anonymous):

I get now! I think I'm over thinking the question. Thank you!